# arXiv:2602.24153v2[math.AG] 02 Mar 2026

### v2

@@ Introduction
In the case d= 3,4d\,{=}\,3,4, the existence of such a vector field implies that either the conclusion of Theorem 2 is satisfied or the curve $C$ is a union of a line and a conic or two conics both intersecting at one point so that 3/d-3/d is a root of the Bernstein-Sato polynomial at a singular point of $C$, see [Sa 25, Remark 8.5]. It is then easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [BaVe 26]) to the following (by using known results in the two variable case [Lo 88], see Remark 3.1 below).
@@
This can be verified rather easily if one admits a formula for the local topological zeta function of a Newton-nondegenerate polynomial of three variables (which is quite amazing), see [DeLo 92]. In the applied case it is possible to deduce this formula from the one for the two variable case (which seems much easier to prove) using the blowup at the origin followed by the ones with line centers together with some computations of rational functions by a computer algebra program such as Macaulay2 or Singular among others. Here it is very surprising that the pole 3/d-3/d disappears  when one calculates $Z^{\rm top}_{f,0}(s)$ even if the Euler number of ${\mathbb{P}}^{2}\,{\setminus}\,C$ does not vanish so that one could expect a counterexample, see Remark 3.2 below.
Note finally that the strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities, see [BaVe 26], where the vanishing of the residue at 3/d-3/d is proved in the needed case. Unfortunately it does not seem to be mentioned there that the following assertion actually holds.
@@
This can be reduced to [BaVe 26, Theorem C] using Proposition 3.3 below, since the strong monodromy conjecture for homogeneous polynomials with isolated singularities is easy. This proposition seems to hold also in the non-reduced case.
@@ 2. Newton-nondegenerate case
We explain the notation only in the case n= 3n\,{=}\,3, see [DeLo 92], [LeVP 11] for the general case. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ or $1$ if two of the coefficients of $\ell_{\sigma}$ vanishes (since $f$ is reduced).
@@ 3. Proof of Theorem 3
In the case $f$ is a homogeneous polynomial, an embedded resolution can be obtained by the blowup of the origin followed by the base change of an embedded resolution of the projective hypersurface by the projection to the exceptional divisor of the first blowup. In view of the definition of local topological zeta function, this implies that 3/d-3/d is the only possible pole of $Z^{\rm top}_{f,0}(s)$ which is not a pole of $Z^{\rm top}_{h,0}(s)$ for a local defining function $h$ of $C$.
@@
The above argument can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the (modified) “global” topological zeta functions of $C$ (which contains the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$) is divisible  by $s+3/d$ so that the pole 3/d-3/d disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the (modified) “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
@@
Proposition 3.3.
Assume a reduced projective hypersurface $Z$ of dimension at least $2$ and degree at least $3$ has only homogeneous isolated singularities which are not ordinary double points, and there is a nonzero linear function $\ell(\nu_{1},\dots,\nu_{n})\,{=}\,\hbox{$\sum$}_{i=1}^{n}\,c_{i}\nu_{i}$ on ${\mathbb{R}}^{n}$ with rational coefficients $c_{i}$ such that ${\rm Supp}_{\bf x}\thinspace f\,{\subset}\,\ell^{-1}(0)$ with $f$ a defining polynomial of $Z$. Then $Z$ is a cone.
@@
We thus get that the $c_{j}$ are independent of $j\,{\neq}\,i$. These must vanish, since $f^{(j)}$ has non-isolated singularities otherwise. Then only $c_{i}$ is nonzero, that is, $f$ is a polynomial of $x_{j}$ for $j\,{\neq}\,i$. This contradicts the assumption that $Z$ is not a cone. So the assertion follows. ∎

### v3

# arXiv:2602.24153v3[math.AG] 03 Mar 2026

@@ Introduction
In the case d= 3,4d\,{=}\,3,4, the existence of such a vector field implies that either the conclusion of Theorem 2 is satisfied or the curve $C$ is a union of a line and a conic or two conics both intersecting at one point so that $\tfrac{3}{d}$ is a root of the Bernstein-Sato polynomial at a singular point of $C$, see [Sa 25, Remark 8.5]. It is then easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [BaVe 26]) to the following (by using known results in the two variable case [Lo 88], see Remark 3.1 below).
@@
This can be verified rather easily if one admits a formula for the local topological zeta function of a Newton-nondegenerate polynomial of three variables (which is quite amazing), see [DeLo 92]. In the applied case it is possible to deduce this formula from the one for the two variable case (which seems much easier to prove) using the blowup at the origin followed by the ones with line centers together with some computations of rational functions by a computer algebra program such as Macaulay2 or Singular among others. Here it is very surprising that the pole $\tfrac{3}{d}$ disappears  when one calculates $Z^{\rm top}_{f,0}(s)$ even if the Euler number of ${\mathbb{P}}^{2}\,{\setminus}\,C$ does not vanish so that one could expect a counterexample, see Remark 3.2 below.
Note finally that the strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities, see [BaVe 26], where the vanishing of the residue at $\tfrac{3}{d}$ is proved in the needed case. Here it does not seem to be mentioned that the following assertion actually holds.
@@
This can be reduced to [BaVe 26, Theorem C] using Lemma 3.3 below, since the strong monodromy conjecture for homogeneous polynomials with isolated singularities is easy. This proposition seems to hold also in the non-reduced case.
@@ 2. Newton-nondegenerate case
We explain the notation in the case n= 3n\,{=}\,3, see [DeLo 92], [LeVP 11] for the general case. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ or $1$ if two of the coefficients of $\ell_{\sigma}$ vanishes (since $f$ is reduced).
@@ 3. Proof of Theorem 3
In the case $f$ is a homogeneous polynomial, an embedded resolution can be obtained by the blowup of the origin followed by the base change of an embedded resolution of the projective hypersurface by the projection to the exceptional divisor of the first blowup. In view of the definition of local topological zeta function, this implies that $\tfrac{3}{d}$ is the only possible pole of $Z^{\rm top}_{f,0}(s)$ which is not a pole of $Z^{\rm top}_{h,0}(s)$ for a local defining function $h$ of $C$.
@@
The above argument can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the (modified) “global” topological zeta functions of $C$ (which contains the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$) is divisible  by $s+\tfrac{3}{d}$ so that the pole $\tfrac{3}{d}$ disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the (modified) “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
@@
Lemma 3.3.
Assume a reduced projective hypersurface $Z$ of dimension at least $2$ and degree at least $3$ has only homogeneous isolated singularities which are not ordinary double points, and moreover there is a nonzero linear function $\ell(\nu_{1},\dots,\nu_{n})\,{=}\,\hbox{$\sum$}_{i=1}^{n}\,c_{i}\nu_{i}$ on ${\mathbb{R}}^{n}$ with rational coefficients $c_{i}$ such that ${\rm Supp}_{\bf x}\thinspace f\,{\subset}\,\ell^{-1}(0)$ with $f$ a defining polynomial of $Z$. Then $Z$ is a cone.
@@
We thus proved that the $c_{j}$ are independent of $j\,{\neq}\,i$. This implies the inclusion
${\rm Supp}_{\bf x}\thinspace f\subset\bigl\{\nu\,{\in}\,{\mathbb{N}}^{n}\mid\hbox{$\sum$}_{j}\,\nu_{j}\,{=}\,d,\,\nu_{i}\,{=}\,c\thinspace\hbox{$\sum$}_{j\,{\neq}\,i}\,\nu_{j}\bigr\}$
for some $c\,{\geqslant}\,0$. Then $f$ is either divisible by $x_{i}$ or a polynomial of $x_{j}$ for $j\,{\neq}\,i$. Here the first case does not occur, since $Z$ has non-isolated singularities otherwise. So the assertion follows. ∎